# Emergent spin order and steady-state superradiance in one-dimensional baths

### v1

@@ III EMERGENT SPIN ORDER
We begin by reviewing a single-mode bad cavity, the simplest setting for atomic synchronization [Meiser09, Meiser10, Xu14, Zhu15]. We then add layers of complexity via one-dimensional reservoirs: a ring cavity with two competing bright jump operators, and a bidirectional waveguide where competition coexists with propagation encoded by coherent interactions.
@@ III.1 Single-mode cavity: superradiant lasing
Figure 2: Order in the single-mode cavity. Wigner function for the electromagnetic field emitted by $N=1024$ atoms coupled to a single-mode cavity, for $\Gamma^{\prime}=20\Gamma_{\rm 1D}$ and $w=N\Gamma_{\rm 1D}/2$. White squares indicate mean-field results.
@@
where $r$ and $\psi$ denote the field amplitude and phase. According to Eq. (16c), phases have an all-to-all homogeneous coupling [Fig. 1(c)] as in the Kuramoto model. In terms of $r$ and $\psi$, Eq. (16c) becomes
@@
For $w$ within the synchronization window this realizes a Kuramoto model with positive coupling, leading to phase synchronization [Zhu15]. In the steady state, atoms emit a field with fixed amplitude $r$ but an arbitrary global phase $\psi$, reflecting the global $U(1)$ symmetry of the master equation (under the transformation $\hat{\sigma}_{+}^{n}\rightarrow e^{i\theta}\hat{\sigma}_{+}^{n}$ $\forall n$). As in conventional single-mode lasers, quantum fluctuations induce phase diffusion of the mean-field solution [Mandel95, Tieri17]. In the complex $E$-plane, the Wigner function forms a ring with radius set by $r$, indicating fixed intensity but undefined phase, as illustrated in Fig. 2. Shot-to-shot the phase is picked by spontaneous symmetry breaking.
@@ III.2 Ring cavity: winner takes all
Figure 3: Order in the ring cavity. Mean-field trajectories and Wigner function in (a) the space of right- and left-field amplitudes, and (b) in the complex plane of the right field. Time is encoded in the line transparency (fainter for ealier times), and each trajectory is colored according to which field dominates in the steady state. In (b), the Wigner function in regions far from the center have been multiplied by a factor of 10 for visualization purposes. Panel (c) shows the magnetization $\mathcal{M}$ for a single realization of the phase dynamics, where the right field prevails. Parameters: $N=1024$, $\Gamma^{\prime}=0$, $w=N\Gamma_{\rm 1D}/4$.
In the ring cavity, the competition between the two modes leads to dynamical spontaneous symmetry breaking of both $U(1)$ (global phase) and mirror (left vs right order) symmetries at the level of individual realizations, as we demonstrate below. In analogy with the single-mode cavity case, we define classical order parameters for left and right emission with amplitudes $r_{L/R}$ and phases $\varphi_{L/R}$ as
$E_{L/R}=\sum_{n=1}^{N}e^{\pm ikz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{L/R}e^{-{\rm i}\varphi_{L/R}}.$ (22)
@@
$\dot{\phi}_{\ell}=\frac{\Gamma_{\rm 1D}s_{\ell}^{z}}{2s^{\perp}_{\ell}}\sum_{\nu=\text{L,R}}r_{\nu}\sin(\varphi_{\nu}-\theta_{\nu,\ell}-\phi_{\ell}),$ (23)
where we have defined the position-dependent phases $\theta_{L/R,n}=\mp kz_{n}$. When pumping produces positive inversion and nonzero coherences, the dynamics involve two competing feedback loops that pull $\phi_{\ell}$ toward either $\phi_{\ell}^{R}=\varphi_{R}-kz_{n}$ or $\phi_{\ell}^{L}=\varphi_{L}+kz_{n}$. Each of these attractors maximizes either $r_{R}$ or $r_{L}$, and the strength of attraction grows with the corresponding field amplitude. Thus, the strongest field at a given time attracts phases toward its fixed point, further amplifying itself and suppressing the rival mode. As captured by Eq. (22) and sketched in Fig. 1(b), the light emitted by each atom into the right/left mode serves as feedback to synchronize the rest of the atoms to the right/left order.
Within the synchronization window, as shown in Fig. 3(a) and (b), mean-field trajectories relax to one of two steady states: either the left field acquires a finite amplitude while the right field is suppressed, or vice versa, reminiscent of a conventional ring-laser behavior [Mandel95]. To visualize the microscopic phase order, we define the “magnetization” for atom $n$ at time $t$ as
@@
where $\Delta\phi_{n}\equiv\phi_{n}-\phi_{n-1}$. The magnetization quantifies whether neighbors are closer to maximizing right ($\mathcal{M}(n,t)\rightarrow 1$) or left ($\mathcal{M}(n,t)\rightarrow-1$) emission. Figure 3(c) shows that, after an initial transient, the phases stabilize to either the right or the left order.
@@
As discussed in Sec. II, the steady state of Eq. (1) is unique. The two solutions therefore do not correspond to distinct steady states of the master equation. Individual decay and incoherent pumping reduce the mirror-inversion symmetry of Eq. (1) for the ring-cavity model to a weak symmetry [Buca12, Albert14] (i.e., a symmetry preserved at the level of the master equation but not at the level of single trajectories). Consistent with this, the two peaks in the Wigner distribution result from stochastic switching between the two dynamical outcomes in measurement-conditioned trajectories [Wiseman09], in contrast to dissipative freezing that occurs in the presence of a strong symmetry [Sanchez19]. As the atom number increases, the Wigner distribution becomes increasingly localized around a single mean-field solution. Consequently, any individual trajectory or experimental realization will remain localized near one of the two mean-field solutions, effectively breaking the symmetry.
The locking to either a global left or right phase order motivates an ansatz for the steady state as an incoherent superposition of spin-coherent states with phases matching the two possible orders [Zhang25, Zhang25B]. We harness the form of the ansatz recently proposed in Ref. [Zhang25B] for transient superradiant dynamics in a waveguide (but here employed to describe the steady state in a ring cavity), which reads
$\hat{\rho}_{\text{ans,RC}}(t_{ss})=\frac{1}{4\pi}\int_{0}^{2\pi}d\varphi\left[\hat{\rho}_{R}(\varphi)+\hat{\rho}_{L}(\varphi)\right],$ (25)
@@
$\begin{aligned} \hat{\rho}_{L/R}(\varphi)&=\bigotimes_{n=1}^{N}\mathinner{|{\theta_{n},\pm kz_{n}+\varphi}\rangle}\mathinner{\langle{\theta_{n},\pm kz_{n}+\varphi}|},\end{aligned}$ (26)
and spin coherent states are defined as $\mathinner{|{\theta,\phi}\rangle}=\cos\frac{\theta}{2}\mathinner{|{e}\rangle}+e^{{\rm i}\phi}\sin\frac{\theta}{2}\mathinner{|{g}\rangle}$. The free angle $\varphi$ in Eq. (25) reflects the global $U(1)$ symmetry. Heterodyne unraveling [Wiseman09] is particularly well-suited to observe the ordering encoded in Eq. (25): although the steady state of Eq. (1) has vanishing atomic coherences due to the $U(1)$ symmetry, measurement backaction generates coherence at the level of individual trajectories. As a result, each realization develops nonzero coherences and, after sufficiently long evolution, tends to synchronize to either the left or the right order.
Observables sensitive to the synchronization direction average to zero in the full density matrix, since roughly half of the realizations lock to each side. Nonetheless, signatures of the order are observable at the level of correlation functions [Zhu15]
@@ III.3 Waveguide: coexistence of left and right order
In addition to the competition between jump operators, the waveguide has a non-zero Hamiltonian that encodes the propagation of light [Goncalves25], dramatically altering the steady state. Heuristically, the distinction from a ring cavity lies in boundary conditions. In a ring, light emitted by any atom into either mode loops around to provide global feedback to the entire ensemble [see the definition of the order parameters in Eq. (22)]. In a waveguide the feedback is directional, as light emitted by atom $\ell$ to the right mode only serves as feedback for atoms located to its right, as sketched in Fig. 1(d). In the mean-field treatment, this asymmetry suggests the definition of right and left position-dependent order parameters
$E_{R}(\ell)$ $=\sum_{z_{n}<z_{\ell}}e^{-{\rm i}kz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{R}(\ell)e^{-{\rm i}\varphi_{R}(\ell)},$ (29a)
$E_{L}(\ell)$ $=\sum_{z_{n}>z_{\ell}}e^{{\rm i}kz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{L}(\ell)e^{-{\rm i}\varphi_{L}(\ell)},$ (29b)
a construction with parallels to earlier approaches that introduced position-dependent order parameters to describe synchronization in systems with inhomogeneous couplings and frustration [Kuramoto02, Abrams04]. In terms of these, the equations for atoms coupled to a waveguide mirror those of the ring cavity, except that the order parameters are now position-dependent. For example, the equation that describes the phase evolution reads
$\dot{\phi}_{\ell}=\frac{\Gamma_{\rm 1D}s_{\ell}^{z}}{s^{\perp}_{\ell}}\sum_{\nu=\text{L,R}}r_{\nu}(\ell)\sin(\varphi_{\nu}(\ell)-\theta_{\nu,\ell}-\phi_{\ell}).$ (30)
In the mean-field treatment, each phase thus evolves under position–dependent feedback. Emission into the right mode accumulates along the array, peaking at the right edge, where atoms are coherent and lock their phases to radiate to the right. Symmetrically, atoms on the left lock to the left order. The impact of coherent interactions is made evident by the magnetization $\mathcal{M}(n,t)$, shown in Fig. 4(a) for a trajectory of the classical dynamics. Unlike the ring cavity [Fig. 3(c)], where a single macroscopic order dominates, atoms in a bidirectional waveguide undergo phase separation between the two orders.
@@
$\hat{\rho}_{\text{ans,WG}}(t_{ss})=\int_{0}^{2\pi}\frac{d\varphi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\varphi_{L}}{2\pi}\sum_{\textbf{A}\in\{R,L\}^{N}}P(\textbf{A})$ (31)
$\times\bigotimes_{n=1}^{N}\mathinner{|{\theta_{n},\phi(A_{n},n,\varphi_{A_{n}})}\rangle}\mathinner{\langle{\theta_{n},\phi(A_{n},n,\varphi_{A_{n}})}|}.$
@@
$\phi(A_{n},n,\varphi_{A_{n}})=\begin{cases}kz_{n}+\varphi_{L}&\text{ if }A_{n}=L\\ -kz_{n}+\varphi_{R}&\text{ if }A_{n}=R.\end{cases}$ (32)
Here, $P(\mathbf{A})$ denotes the probability of a given configuration of local orderings. For example, $\mathbf{A}=[L,L,R,L\ldots]$ means that the first two atoms align with the left order, the third aligns with the right order, and so on. The independent integration over $\varphi_{R}$ and $\varphi_{L}$ reflects the assumption that the relative phase between the left and right ordered domains is completely undetermined in the steady state (at least, on average). It also respects the global $U(1)$ symmetry.
@@
  • Interdomain correlations [(IV) and (V)]: These correspond to atom pairs located in opposite domains. For a fixed relative phase between orders, each pair exhibits a well-defined phase relation, $\arg[C_{nm}]\simeq\pm[k(z_{n}+z_{m})+\varphi_{L}-\varphi_{R}]$. However, averaging over $\varphi_{R}$ and $\varphi_{L}$ washes out this phase relation, restoring a cosine-like pattern.
@@
We note that the ansatz in Eq. (31) is a phenomenological model intended to reproduce the behavior of correlations after averaging over trajectories. A more detailed analysis of individual trajectories may provide further insight on how quantum fluctuations influence the mean-field order. For example, although the relative phase between the left and right order parameters is assumed to be arbitrary, more general models could be developed by introducing a probability distribution of the relative phase, $P(\varphi_{R}-\varphi_{L})$. The present ansatz reproduces the correct average correlations provided that the probability distribution of this relative phase is an even function. A possible refinement of the model would be to allow for atoms that are not synchronized with either order parameter. Nonetheless, the minimal model in Eq. (31) already demonstrates that the mean-field intuition is consistent with the TWA numerical results.
@@ IV PROPERTIES OF THE EMITTED LIGHT
Here we determine the pumping thresholds for synchronization and superradiant emission, and show that the emitted intensity scales as $N^{2}$ for all three spin models. Using an extension of the TWA to compute multi-time correlators [Guardiola25] together with an $SU(4)$ representation of the Liouvillian for the single-mode and ring cavities [Xu15, Lee25], we then analyze the degree of first-order coherence of the field by studying the spectral linewidth.

### v2

@@ III EMERGENT SPIN ORDER
We begin by reviewing a single-mode bad cavity, the simplest setting for atomic synchronization [42, 41, 66, 76]. We then add layers of complexity via one-dimensional reservoirs: a ring cavity with two competing bright jump operators, and a bidirectional waveguide where competition coexists with propagation encoded by coherent interactions.
@@ III.1 Single-mode cavity: superradiant lasing
Figure 2: Order in the single-mode cavity. Wigner function for the electromagnetic field emitted by $N=1024$ atoms coupled to a single-mode cavity, for $\Gamma^{\prime}=20\Gamma_{\rm 1D}$ and $w=N\Gamma_{\rm 1D}/2$. White squares indicate examples of mean-field results.
@@
where $r$ and $\psi$ denote the field amplitude and phase. According to Eq. (16c), phases have an all-to-all homogeneous coupling [Fig. 1(a)] as in the Kuramoto model. In terms of $r$ and $\psi$, Eq. (16c) becomes
@@
For $w$ within the synchronization window this realizes a Kuramoto model with positive coupling, leading to phase synchronization [76]. In the steady state, atoms emit a field with fixed amplitude $r$ but an arbitrary global phase $\psi$, reflecting the global $U(1)$ symmetry of the master equation. As in conventional single-mode lasers, quantum fluctuations induce phase diffusion of the mean-field solution [38, 61]. In the complex $E$-plane, the Wigner function forms a ring with radius set by $r$, indicating fixed intensity but undefined phase, as illustrated in Fig. 2. Shot-to-shot the phase is picked by spontaneous symmetry breaking.
@@ III.2 Ring cavity: winner takes all
Figure 3: Order in the ring cavity. Mean-field trajectories and Wigner function in (a) the space of right- and left-field amplitudes, and (b) in the complex plane of the right field. Diamonds indicate the final configurations of the mean-field dynamics. Time is encoded in the line transparency (fainter for earlier times), and each trajectory is colored according to which field dominates in the steady state. In (b), the Wigner function in regions far from the center have been multiplied by a factor of 10 for visualization purposes. Panel (c) shows the magnetization $\mathcal{M}$ for a single realization of the phase dynamics, where the right field prevails. Parameters: $N=1024$, $\Gamma^{\prime}=0$, $w=N\Gamma_{\rm 1D}/4$.
In the ring cavity, the competition between the two modes leads to dynamical spontaneous symmetry breaking of both $U(1)$ (global phase) and mirror (left vs right order) symmetries at the level of individual realizations, as we demonstrate below. In analogy with the single-mode cavity case, we define classical order parameters for left and right emission with amplitudes $r_{L/R}$ and phases $\psi_{L/R}$ as
$E_{L/R}=\sum_{n=1}^{N}e^{\pm ikz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{L/R}e^{-{\rm i}\psi_{L/R}}.$ (22)
@@
$\dot{\phi}_{\ell}=\frac{\Gamma_{\rm 1D}s_{\ell}^{z}}{2s^{\perp}_{\ell}}\sum_{\nu=\text{L,R}}r_{\nu}\sin(\psi_{\nu}-\theta_{\nu,\ell}-\phi_{\ell}),$ (23)
where we have defined the position-dependent phases $\theta_{L/R,n}=\mp kz_{n}$. When pumping produces positive inversion and nonzero coherences, the dynamics involve two competing feedback loops that pull $\phi_{\ell}$ toward either $\phi_{\ell}^{R}=\psi_{R}-kz_{n}$ or $\phi_{\ell}^{L}=\psi_{L}+kz_{n}$. Each of these attractors maximizes either $r_{R}$ or $r_{L}$, and the strength of attraction grows with the corresponding field amplitude. Thus, the strongest field at a given time attracts phases toward its fixed point, further amplifying itself and suppressing the rival mode. As captured by Eq. (22) and sketched in Fig. 1(b), the light emitted by each atom into the right/left mode serves as feedback to synchronize the rest of the atoms to the right/left order.
Within the synchronization window, as shown in Fig. 3(a) and (b), mean-field trajectories relax to one of two steady states: either the left field acquires a finite amplitude while the right field is suppressed, or vice versa, reminiscent of a conventional ring-laser behavior [38]. To visualize the microscopic phase order, we define the “magnetization” for atom $n$ at time $t$ as
@@
where $\Delta\phi_{n}\equiv\phi_{n}-\phi_{n-1}$. The magnetization quantifies whether neighbors are closer to maximizing right ($\mathcal{M}(n,t)\rightarrow 1$) or left ($\mathcal{M}(n,t)\rightarrow-1$) emission. Figure 3(c) shows that, after an initial transient, all the phases stabilize to either the right or the left order.
@@
As discussed in Sec. II, the steady state of Eq. (1) is unique. The two solutions therefore do not correspond to distinct steady states of the master equation. Individual decay and incoherent pumping reduce the mirror-inversion symmetry of Eq. (1) for the ring-cavity model to a weak symmetry [8, 3] (i.e., a symmetry preserved at the level of the master equation but not at the level of single trajectories). Consistent with this, the two peaks in the Wigner distribution result from stochastic switching between the two dynamical outcomes in measurement-conditioned trajectories [64], in contrast to dissipative freezing that occurs in the presence of a strong symmetry [54]. As the atom number increases, the Wigner distribution becomes increasingly localized around a single mean-field solution. Consequently, any individual trajectory or experimental realization will remain localized near one of the two mean-field solutions, effectively breaking the symmetry.
The locking to either a global left or right phase order motivates an ansatz for the steady state as an incoherent superposition of spin-coherent states with phases matching the two possible orders [72, 71]. We harness the form of the ansatz recently proposed in Ref. [71] for transient superradiant dynamics in a waveguide (but here employed to describe the steady state in a ring cavity), which reads
$\hat{\rho}_{\text{ans,RC}}(t_{ss})=\frac{1}{4\pi}\int_{0}^{2\pi}d\psi\left[\hat{\rho}_{R}(\psi)+\hat{\rho}_{L}(\psi)\right],$ (25)
@@
$\begin{aligned} \hat{\rho}_{L/R}(\psi)&=\bigotimes_{n=1}^{N}\mathinner{|{\theta_{n},\pm kz_{n}+\psi}\rangle}\mathinner{\langle{\theta_{n},\pm kz_{n}+\psi}|},\end{aligned}$ (26)
and spin coherent states are defined as $\mathinner{|{\theta,\phi}\rangle}=\cos\frac{\theta}{2}\mathinner{|{e}\rangle}+e^{{\rm i}\phi}\sin\frac{\theta}{2}\mathinner{|{g}\rangle}$. The free angle $\psi$ in Eq. (25) reflects the global $U(1)$ symmetry. Heterodyne unraveling [64] is particularly well-suited to observe the ordering encoded in Eq. (25): although the steady state of Eq. (1) has vanishing atomic coherences due to the $U(1)$ symmetry, measurement backaction generates coherence at the level of individual trajectories. As a result, each realization develops nonzero coherences and, after sufficiently long evolution, tends to synchronize to either the left or the right order.
Observables sensitive to the synchronization direction average to zero in the full density matrix, since roughly half of the realizations lock to each side. Nonetheless, signatures of the order are observable at the level of correlation functions [76]
@@ III.3 Waveguide: coexistence of left and right order
In addition to the competition between jump operators, the waveguide has a non-zero Hamiltonian that encodes the propagation of light [23], dramatically altering the steady state. Heuristically, the distinction from a ring cavity lies in boundary conditions. In a ring, light emitted by any atom into either mode loops around to provide global feedback to the entire ensemble [see the definition of the order parameters in Eq. (22)]. In a waveguide the feedback is directional, as light emitted by atom $\ell$ to the right mode only serves as feedback for atoms located to its right, as sketched in Fig. 1(c). In the mean-field treatment, this asymmetry suggests the definition of right and left position-dependent order parameters
$E_{R}(\ell)$ $=\sum_{z_{n}<z_{\ell}}e^{-{\rm i}kz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{R}(\ell)e^{-{\rm i}\psi_{R}(\ell)},$ (29a)
$E_{L}(\ell)$ $=\sum_{z_{n}>z_{\ell}}e^{{\rm i}kz_{n}}s^{\perp}_{n}e^{-{\rm i}\phi_{n}}\equiv r_{L}(\ell)e^{-{\rm i}\psi_{L}(\ell)},$ (29b)
a construction with parallels to earlier approaches that introduced position-dependent order parameters to describe synchronization in systems with inhomogeneous couplings and frustration [31, 1]. In terms of these, the equations for atoms coupled to a waveguide mirror those of the ring cavity, except that the order parameters are now position-dependent. For example, the equation that describes the phase evolution reads
$\dot{\phi}_{\ell}=\frac{\Gamma_{\rm 1D}s_{\ell}^{z}}{s^{\perp}_{\ell}}\sum_{\nu=\text{L,R}}r_{\nu}(\ell)\sin(\psi_{\nu}(\ell)-\theta_{\nu,\ell}-\phi_{\ell}).$ (30)
In the mean-field treatment, each phase thus evolves under position–dependent feedback. Emission into the right mode accumulates along the array, peaking at the right edge, where atoms are coherent and lock their phases to radiate to the right. Symmetrically, atoms on the left lock to the left order. The impact of coherent interactions is made evident by the magnetization $\mathcal{M}(n,t)$, shown in Fig. 4(a) for a trajectory of the classical dynamics. Unlike the ring cavity [Fig. 3(c)], where a single macroscopic order dominates, atoms in a bidirectional waveguide undergo phase separation between the two orders. Most realizations of the mean field dynamics thus reach a steady state where the right and left field intensity are approximately identical.
@@
$\hat{\rho}_{\text{ans,WG}}(t_{ss})=\int_{0}^{2\pi}\frac{d\psi_{R}}{2\pi}\int_{0}^{2\pi}\frac{d\psi_{L}}{2\pi}\sum_{\textbf{A}\in\{R,L\}^{N}}P(\textbf{A})$ (31)
$\times\bigotimes_{n=1}^{N}\mathinner{|{\theta_{n},\phi(A_{n},n,\psi_{A_{n}})}\rangle}\mathinner{\langle{\theta_{n},\phi(A_{n},n,\psi_{A_{n}})}|},$
@@
$\phi(A_{n},n,\psi_{A_{n}})=\begin{cases}kz_{n}+\psi_{L}&\text{ if }A_{n}=L\\ -kz_{n}+\psi_{R}&\text{ if }A_{n}=R.\end{cases}$ (32)
Here, $P(\mathbf{A})$ denotes the probability of a given configuration of local orderings. For example, $\mathbf{A}=[L,L,R,L\ldots]$ means that the first two atoms align with the left order, the third aligns with the right order, and so on. The independent integration over $\psi_{R}$ and $\psi_{L}$ reflects the assumption that the relative phase between the left and right ordered domains is completely undetermined in the steady state (at least, on average). It also respects the global $U(1)$ symmetry.
@@
  • Interdomain correlations [(IV) and (V)]: These correspond to atom pairs located in opposite domains. For a fixed relative phase between orders, each pair exhibits a well-defined phase relation, $\arg[C_{nm}]\simeq\pm[k(z_{n}+z_{m})+\psi_{L}-\psi_{R}]$. However, averaging over $\psi_{R}$ and $\psi_{L}$ washes out this phase relation, restoring a cosine-like pattern.
@@
We note that the ansatz in Eq. (31) is a phenomenological model intended to reproduce the behavior of correlations after averaging over trajectories. A more detailed analysis of individual trajectories may provide further insight on how quantum fluctuations influence the mean-field order. For example, although the relative phase between the left and right order parameters is assumed to be arbitrary, more general models could be developed by introducing a probability distribution of the relative phase, $P(\psi_{R}-\psi_{L})$. The present ansatz reproduces the correct average correlations provided that the probability distribution of this relative phase is an even function. A possible refinement of the model would be to allow for atoms that are not synchronized with either order parameter. Nonetheless, the minimal model in Eq. (31) already demonstrates that the mean-field intuition is consistent with the TWA numerical results.
@@ IV PROPERTIES OF THE EMITTED LIGHT
Here we determine the pumping thresholds for synchronization and superradiant emission, and show that the emitted intensity scales as $N^{2}$ for all three spin models. Using an extension of the TWA to compute multi-time correlators [26] together with an $SU(4)$ representation of the Liouvillian for the single-mode cavity [65], we then analyze the degree of first-order coherence of the field by studying the spectral linewidth.